OpenStudy (anonymous):

Let f(x) = 5x + 12. Find f-1(x).

OpenStudy (anonymous):

\[ x=5f^{-1}(x)+12 \]Now solve for \(f^{-1}(x)\)

OpenStudy (anonymous):

@wio I have no idea how, but thankyouu !! I appreciate it

OpenStudy (anonymous):

What math are you in?

OpenStudy (anonymous):

Alg.2 , and math is my WORST subject

OpenStudy (anonymous):

Okay so start buy subtracting both sides by \(12\).

OpenStudy (anonymous):

\[ x-12=5f^{-1}(x) \]

OpenStudy (anonymous):

What do you think is the next step?

OpenStudy (anonymous):

Divide by 5 ?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

and from there I get the final answer?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

so it would be x−12/5=f−1(x) ?

OpenStudy (anonymous):

\[ f^{-1}(x) = \frac x5-\frac{12}5 \]

OpenStudy (anonymous):

Thankyouuu ! :D

OpenStudy (anonymous):

Remember that \[ f(f^{-1}(x)) = x \]

OpenStudy (anonymous):

I wrote it down, thankyou ! :] @wio

OpenStudy (anonymous):

Would the answer 1 over 5x+12 ? @wio

OpenStudy (anonymous):

*be

OpenStudy (anonymous):

We already got the answer.

OpenStudy (anonymous):

x-12 over 5 Got it, thankyouu